OpenStudy (anonymous):

A bank offers a CD that pays a simple interest rate of 5.5%. How much must you put in this cd now in order to have $8000 for a kitchen remodeling project in two years? Could you please show work?

OpenStudy (anonymous):

8000=C(1+2x5.5%) This gives C=7207.2

OpenStudy (mertsj):

\[A=P+Prt\] 8000=P(1+rt) 8000=P(1+.055(2)) 8000=P(1.11) P=8000/1.11=7207.20

OpenStudy (anonymous):

Thanks for your help, I appreciate it